OpenStudy (jjuden):

Find the GCF of the terms of the polynomial. 16x^3 + 28x^5y

OpenStudy (jjuden):

A 112x^5y B 4x^5y C 4x^3 D 8x^5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

What does GCF stand for?

OpenStudy (skullpatrol):

$$\Huge 16x^3 + 28x^5y$$

OpenStudy (skullpatrol):

What is the GCF of 16 and 28 @jjuden?

OpenStudy (skullpatrol):

Do you know what a "factor" is @jjuden?

OpenStudy (jjuden):

no

OpenStudy (skullpatrol):

When two numbers are multiplied each of the numbers is a "factor" of the product.

OpenStudy (skullpatrol):

16 is the product.

OpenStudy (skullpatrol):

what numbers multiplied together give you 16?

OpenStudy (skullpatrol):

1*16=16 2*8=16 4*4=16

OpenStudy (skullpatrol):

So 1, 2, 4, 8, and 16 are all "factors" of 16, right @jjuden?

OpenStudy (skullpatrol):

Are you still there @jjuden?

OpenStudy (jjuden):

yes

OpenStudy (skullpatrol):

Can you do the same for 28?

OpenStudy (skullpatrol):

$$16x^3 + 28x^5y$$

OpenStudy (skullpatrol):

Yes or no @jjuden?

OpenStudy (skullpatrol):

Follow the same steps I gave for 16, that is how you learn math :-)